# Chinese Academy of Sciences, Beijing 100190, China

### v1

@@ 3.3 Numerical results of neutrino decoupling
Figure 4: The spectral distortions of neutrinos and antineutrinos in the flavor basis (top left) and mass basis (lower left) with respect to the Fermi-Dirac distribution with corresponding $\xi_{\nu}$. The meaning of the curves is the same as in the previous plot Fig. 3, while the spectral distortions of tauon neutrino in the left panel are not shown for brevity. The zoomed version of the spectral distortions in the flavor and mass basis are shown in the right panel.
@@
Considering that the density distributions of neutrinos and antineutrinos coincide with the corresponding thermal Fermi-Dirac distributions for small momentum, it is natural to define the spectral distortions as $\delta\varrho_{\nu,\bar{\nu}}\equiv\delta\varrho_{\nu,\bar{\nu}}/f_{FD}(\pm\xi)-1$, where the plus sign corresponds to neutrinos and the minus sign to antineutrinos. The resulting spectral distortions as a function of the momentum $y$ in the flavor and mass basis are presented in the upper left and upper right panels of Fig. 4, respectively. The meaning of the curves is the same as in the previous plot Fig. 3, while the spectral distortions of the tauon neutrinos are not shown in the left panels, since they almost coincide with the spectral distortions of the muon neutrinos. We also plot the zoomed version of the spectral distortions in the lower panels. Interestingly, the spectral distortions for neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ from each other and lie on different sides of the spectral distortions with $\xi_{\nu}=0$. The antineutrinos with non-zero $\xi_{\nu}$ possess a relatively larger spectral distortion than that with $\xi_{\nu}=0$, while the neutrinos possess a relatively smaller spectral distortion. In addition, the larger the $\xi_{\nu}$, the larger the discrepancy between the spectral distortions of neutrinos and antineutrinos, clearly seen from the zoomed plots. Such properties are both valid in the flavor basis and in the mass basis. Finally, the above conclusions about neutrino density matrices also hold for negative $\xi_{\nu}$, except that the density matrices for neutrinos and antineutrinos with $-\xi_{\nu}$ are exchanged comparing with those for $\xi_{\nu}$. In other words, the density distributions and spectral distortions for antineutrinos in Fig. 3 and Fig. 4 are also the density distributions and spectral distortions for neutrinos with the opposite $\xi_{\nu}$.
@@
As a brief conclusion of this section, we have presented the most accurate results of neutrino decoupling in the presence of primordial neutrino asymmetries and discussed the resulting $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos. In the next sections, we will study the effects of primordial neutrino asymmetries on the processes of BBN, CMB and structure formation, while emphasizing the corrections induced by the spectral distortions of neutrinos and antineutrinos.
@@ 4 Impacts of the Primordial Neutrino Asymmetry on BBN
Soon after neutrino decoupling, there is an epoch of primordial nucleosynthesis or the so-called Big Bang Nucleosynthesis (BBN), where the light elements are produced after protons and neutrons decouple from the SM thermal bath. Since the decoupling of protons depends on electron neutrinos, the resulting abundances of light elements after BBN depends on the properties of electron neutrinos and antineutrinos. BBN is affected by primordial neutrino asymmetries in three ways as follow.
@@
    The leading effect is the change in the weak interaction rate at the onset of BBN due to the chemical potential of the electron neutrinos, since the conversion between proton and neutron is related to the electron neutrino through the weak interaction, so that the difference of the chemical potentials of proton and neutron is decided by the chemical potential of electron neutrino, $i.e.,$ $\mu_{n}-\mu_{p}=\mu_{e}-\mu_{\nu_{e}}\simeq-\mu_{\nu_{e}}$. Therefore, the resulting neutron-to-proton ratio and the subsequent BBN processes are modified according to the absolute value of the electron neutrino chemical potential and its sign.
@@
    In addition, the expansion rate of the universe during BBN is changed by the additional neutrino energy density contributed by the chemical potentials. As we have seen in the last section, the neutrino energy density after neutrino decoupling is usually represented by $N_{\rm eff}$, and we have $N_{\rm eff}\simeq 3.044+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$. The corresponding correction in BBN therefore depends on the absolute asymmetries of all three neutrino flavors.
@@
    Beyond the approximation of the thermal Fermi-Dirac distributions for neutrinos and antineutrinos, the spectral distortions resulting from the non-instantaneous decoupling of the neutrinos will also induce a correction of the weak interaction rate and thus affect the BBN. Such corrections are usually neglected in previous studies, especially in the presence of primordial neutrino asymmetries.
In the following, we first review the general treatment of the weak interaction rate and introduce the new correction induced by the neutrino spectral distortions, then we discuss the resulting light element abundances from the BBN.
@@ 4.1 The weak correction from neutrino spectral distortions
At the onset of the BBN, the conversion between neutron ($n$) and proton ($p$) is dominated by following interaction reactions,
@@
In the remainder of this section, we will omit the subscript for the electron-neutrino and the superscript for the electron/positron as long as there is no ambiguity. Due to the forceful coupling between photons and electrons and positrons, it is appropriate to assume that all particles expect for neutrinos are in local thermal equilibrium. Therefore, the neutron and proton number densities obey the Boltzmann equations
@@
$\dot{n}_{p}+3Hn_{p}$ $=-n_{p}\Gamma_{p\to n}+n_{n}\Gamma_{n\to p},$ (69)
where the reaction rates $n_{n}\Gamma_{n\to p}$ and $n_{p}\Gamma_{p\to n}$ incorporate all relevant collision terms
@@
And the reaction rate for a specific interaction reaction with the nucleon $a$ in the initial state and nucleon $b$ in the final state takes the form as
@@
where $\mathrm{d}\pi_{i}=\frac{1}{(2\pi)^{3}}\frac{\mathrm{d}^{3}\mathbf{p}_{i}}{2E_{i}}$, $\alpha_{e}(\alpha_{\nu})=1$ if the electron (neutrino) is in the initial state and $\alpha_{e}(\alpha_{\nu})=-1$ if the electron (neutrino) is in the final state. In addition, $f_{i}(E)\equiv\frac{1}{e^{E/T_{\gamma}}+1}$ for $i=a,b,e$ is the Fermi-Dirac distribution with vanishing chemical potential, and we have used the relation $1-f_{i}(E)=f_{i}(-E)$ to incorporate the corresponding Pauli blocking factors. Lastly, $\tilde{f}_{\nu}(E_{\nu})$ equals $\varrho_{\nu}(E_{\nu})\,(\varrho_{\bar{\nu}}(E_{\nu}))$ for the neutrino (antineutrino) in the initial state and equals $1-\varrho_{\nu}(E_{\nu})\,(1-\varrho_{\bar{\nu}}(E_{\nu}))$ for the neutrino (antineutrino) in the initial state.
@@
with $E=\sqrt{p^{2}+m_{e}^{2}}$ is the energy of the electron and $K=1/(\tau_{n}\lambda_{0}m_{e}^{5})$, where $\tau_{n}$ is the life time of the neutron and $\lambda_{0}$ the Born coefficient. The $\chi^{\pm}$ function is defined as
$\chi_{\pm}(E)\equiv(E_{\nu}^{\mp})^{2}f_{e}(-E)f_{\nu}^{\pm}(E^{\mp}_{\nu})\,,\,\,E^{\mp}_{\nu}\equiv E\mp\Delta,$ (75)
@@
where $\Delta\equiv m_{n}-m_{p}$ is the energy gap between neutron and proton. The first term in Eq. (73) corresponds to the $n\to p+e$ processes Eq. (65) or Eq. (66) depending on whether $E-\Delta$ is positive or negative, and the second term corresponds to the $n+e\to p$ process Eq. (67) and similarly for the terms in Eq. (74). Note that here $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ for the processes where the neutrino is in the initial state, or its opposite $-E_{\nu}$ for the processes where the neutrino is in the final state. And the term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ gives the corresponding distribution or the Pauli-blocking factor by the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
@@
Now we introduce the corrections induced by the spectral distortions of neutrinos and antineutrinos. In the infinite nucleon mass limit, the corrections to the Born rates can be written as
$\delta\bar{\Gamma}_{n\to p}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\delta\chi_{+}(E)+\delta\chi_{+}(-E)]\,,$ (76)
$\delta\bar{\Gamma}_{p\to n}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\delta\chi_{-}(E)+\delta\chi_{-}(-E)]\,,$ (77)
with
$\delta\chi_{\pm}(E)\equiv$ $(E_{\nu}^{\mp})^{2}f_{e}(-E){\rm sgn}(E^{\mp}_{\nu}){}$ (78)
×$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,,$
@@
where $\Theta(x)$ is the Heaviside step function, $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ are the spectral distortions of electron neutrinos and antineutrinos. For a particular process, the various sgn functions and the $\Theta$ functions together represent the nature and the position of the neutrino particle: Whether it is a neutrino or an antineutrino, and whether it is in the initial state or the final state. We also note that $T_{\rm CM}$ and $T_{\nu}$ has slight difference after neutrino decoupling. If we neglect such difference and set $\xi_{\nu}=0$, so that $\delta\varrho_{\nu}=\delta\varrho_{\bar{\nu}}$, then Eq. (78) coincides with Eq. (29) in Ref. Froustey:2019owm , where they considered the corrections to the weak rates from neutrino spectral distortions in the SM.
@@
Finally, we briefly review the other corrections to the weak interaction rates that we have considered in this work. First are the zero temperature radiative corrections, including the so-called pure radiative corrections corresponding to the emission and absorption of a virtual photon in the interaction and the so-called bremsstrahlung correction due to the emission of a real photon. Second are the finite temperature radiative corrections due to the finite temperature corrections to the proton-electron interactions and the electron energy. Third and fourth are the corrections from the finite nucleon mass and the weak magnetism. Combining all these corrections, the total weak reaction rates are given by
@@
where the superscript ’RC0’, ’SD’, ’T’, ’BS’ and ’FM’ represent the zero temperature radiative correction, the spectral distortion, the finite temperature radiative correction, the bremsstrahlung and the finite nucleon mass correction, respectively. And the form for the $p\to n$ rates is similar. Note that the effects of different corrections may be coupled. For example, the effects of the zero temperature radiative correction on the corrections induced by the spectral distortions are also considered in this work. Specifically, the Born rates with the zero temperature radiative correction take the form of
${\Gamma}^{{\rm RC}0}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\mathcal{F}_{+}(E)R(E,|\Delta-E|)\chi_{+}(E)+\mathcal{F}_{+}(-E)R(E,\Delta+E)\chi_{+}(-E)]\,,$ (80)
${\Gamma}^{{\rm RC}0}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\mathcal{F}_{-}(E)R(E,\Delta+E)\chi_{-}(E)+\mathcal{F}_{-}(-E)R(E,|E-\Delta|)\chi_{-}(-E)]\,.$ (81)
The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in $e.g.,$   Ref. Pitrou:2018cgg and are not repeated here for brevity. We see that the zero temperature radiative correction only modifies the coefficients of the $\chi_{\pm}$ function. The $\delta\Gamma^{\rm RC0,SD}_{n\to p}$ and $\delta\Gamma^{\rm RC0,SD}_{p\to n}$ can also be obtained simply by the replacement of $\chi_{\pm}\to\delta\chi_{\pm}$. In principle, the other corrections are also affected by the spectral distortions of neutrinos and antineutrinos. However, such effects must be negligible since the corrections themselves are already sub-leading. Therefore, we neglect these interferences in this work.

### v2

@@ 3.3 Numerical Results of Neutrino Decoupling
Figure 4: Spectral distortions of neutrinos and antineutrinos in the flavor basis (left panel) and mass basis (right panel), relative to the Fermi-Dirac distribution with the corresponding $\xi_{\nu}$. The curves follow the same convention as in Fig. 3, but the spectral distortions of tau neutrinos in the left panel are omitted for brevity.
Since the density distributions of neutrinos and antineutrinos coincide with the thermal Fermi-Dirac distributions at small momenta, it is natural to define the spectral distortions as
$\delta\varrho_{\nu,\bar{\nu}}\equiv\frac{\delta\varrho_{\nu,\bar{\nu}}}{f_{FD}(\pm\xi)}-1\,,$ (65)
@@
where the plus sign corresponds to neutrinos and the minus sign to antineutrinos. The resulting spectral distortions as a function of momentum $y$ in the flavor and mass bases are shown in the left and right panels of Fig. 4, respectively. The curves follow the same convention as in Fig. 3, with tau neutrino spectral distortions omitted in the left panels, as they nearly coincide with those of muon neutrinos. Notably, the spectral distortions for neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ and lie on opposite sides of the spectral distortions for $\xi_{\nu}=0$. Antineutrinos with non-zero $\xi_{\nu}$ exhibit larger spectral distortions compared to those for $\xi_{\nu}=0$, while neutrinos display smaller distortions. Moreover, the larger the $\xi_{\nu}$, the greater the discrepancy between the spectral distortions of neutrinos and antineutrinos. These properties hold in both the flavor and mass bases and arise from the differing Pauli-blocking effects for neutrinos and antineutrinos in the presence of primordial neutrino asymmetries. Finally, the above conclusions regarding neutrino density matrices also apply to negative $\xi_{\nu}$, with the density matrices for neutrinos and antineutrinos swapped compared to those for positive $\xi_{\nu}$. In other words, the density distributions and spectral distortions of antineutrinos in Figs. 3 and 4 also apply to neutrinos with opposite $\xi_{\nu}$.
@@
In summary, we have presented the most accurate results of neutrino decoupling in the presence of primordial neutrino asymmetries and discussed the corresponding $N_{\rm eff}$ and spectral distortions of neutrinos and antineutrinos. In the following sections, we will explore the effects of primordial neutrino asymmetries on BBN, CMB, and structure formation, focusing on the corrections induced by the spectral distortions of neutrinos and antineutrinos.
@@ 4 Impacts of the Primordial Neutrino Asymmetry on BBN
Soon after neutrino decoupling, primordial nucleosynthesis, or Big Bang Nucleosynthesis (BBN), occurs, leading to the production of light elements as protons and neutrons decouple from the Standard Model (SM) thermal bath. The decoupling of protons is influenced by electron neutrinos, causing the resulting abundances of light elements after BBN to depend on the properties of both electron neutrinos and antineutrinos. BBN is impacted by primordial neutrino asymmetries in three ways (for reviews, see, $e.g.,$ Pitrou:2018cgg ):
@@
    The primary effect is the alteration of the weak interaction rate at the onset of BBN, which is influenced by the chemical potential of electron neutrinos. The conversion between protons and neutrons is mediated by the weak interaction, leading to a difference in the chemical potentials of protons and neutrons determined by the chemical potential of electron neutrinos, i.e., $\mu_{n}-\mu_{p}=\mu_{e}-\mu_{\nu_{e}}\simeq-\mu_{\nu_{e}}$. Consequently, the resulting neutron-to-proton ratio and subsequent BBN processes are modified according to the absolute value and sign of the electron neutrino chemical potential.
@@
    Additionally, the expansion rate of the universe during BBN is affected by the increased neutrino energy density contributed by the chemical potentials. As discussed in the previous section, the neutrino energy density following neutrino decoupling is typically represented by $N_{\rm eff}$, which can be expressed as $N_{\rm eff}\simeq 3.044+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$. Thus, the corresponding corrections in BBN depend on the absolute asymmetries of all three neutrino flavors.
@@
    Beyond the approximation of thermal Fermi-Dirac distributions for neutrinos and antineutrinos, spectral distortions arising from the non-instantaneous decoupling of neutrinos will also induce corrections to the weak interaction rate, thereby affecting BBN. Such corrections have often been overlooked in previous studies, particularly in the context of primordial neutrino asymmetries.
In the following sections, we will first review the general treatment of the weak interaction rate and introduce the new corrections induced by neutrino spectral distortions, after which we will discuss the resulting light element abundances from BBN.
@@ 4.1 The Weak Correction from Neutrino Spectral Distortions
At the onset of BBN, the conversion between neutron ($n$) and proton ($p$) is dominated by the following interaction reactions:
@@
In the remainder of this section, we will omit the subscript for the electron neutrino and the superscript for the electron/positron whenever there is no ambiguity. Due to the strong coupling between photons, electrons, and positrons, it is appropriate to assume that all particles except for neutrinos are in local thermal equilibrium. Consequently, the number densities of neutrons and protons obey the Boltzmann equations:
@@
$\dot{n}_{p}+3Hn_{p}$ $=-n_{p}\Gamma_{p\to n}+n_{n}\Gamma_{n\to p}.$ (70)
Here, the reaction rates $n_{n}\Gamma_{n\to p}$ and $n_{p}\Gamma_{p\to n}$ incorporate all relevant collision terms:
@@
The reaction rate for a specific interaction involving nucleon $a$ in the initial state and nucleon $b$ in the final state is given by:
@@
where $\mathrm{d}\pi_{i}=\frac{1}{(2\pi)^{3}}\frac{\mathrm{d}^{3}\mathbf{p}_{i}}{2E_{i}}$, $\alpha_{e}(\alpha_{\nu})=1$ if the electron (neutrino) is in the initial state and $\alpha_{e}(\alpha_{\nu})=-1$ if it is in the final state. Additionally, $f_{i}(E)\equiv\frac{1}{e^{E/T_{\gamma}}+1}$ for $i=a,b,e$ is the Fermi-Dirac distribution with a vanishing chemical potential, and we have used the relation $1-f_{i}(E)=f_{i}(-E)$ to account for the corresponding Pauli blocking factors. Lastly, $\tilde{f}_{\nu}(E_{\nu})$ equals $\varrho_{\nu}(E_{\nu})$ (or $\varrho_{\bar{\nu}}(E_{\nu})$) for the neutrino (or antineutrino) in the initial state, and equals $1-\varrho_{\nu}(E_{\nu})$ (or $1-\varrho_{\bar{\nu}}(E_{\nu})$) for those in the final state.
@@
where $E=\sqrt{p^{2}+m_{e}^{2}}$ represents the energy of the electron and $K=1/(\tau_{n}\lambda_{0}m_{e}^{5})$, with $\tau_{n}$ being the lifetime of the neutron and $\lambda_{0}$ the Born coefficient. The $\chi^{\pm}$ functions are defined as
$\chi_{\pm}(E)\equiv(E_{\nu}^{\mp})^{2}f_{e}(-E)f_{\nu}^{\pm}(E^{\mp}_{\nu})\,,\quad E^{\mp}_{\nu}\equiv E\mp\Delta,$ (76)
@@
where $\Delta\equiv m_{n}-m_{p}$ is the energy gap between the neutron and proton. The first term in (74) corresponds to the $n\to p+e$ processes (66) or (67), depending on whether $E-\Delta$ is positive or negative, while the second term corresponds to the $n+e\to p$ process (68). Similarly, the terms in (75) can be interpreted. Here, $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ when the neutrino is in the initial state, and $-E_{\nu}$ when it is in the final state. The term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ provides the corresponding distribution or Pauli-blocking factor, following the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
@@
Now, we introduce the corrections induced by the spectral distortions of neutrinos and antineutrinos. In the infinite nucleon mass limit, the corrections to the Born rates can be expressed as follows:
$\delta\bar{\Gamma}_{n\to p}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\delta\chi_{+}(E)+\delta\chi_{+}(-E)\right]\,,$ (77)
$\delta\bar{\Gamma}_{p\to n}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\delta\chi_{-}(E)+\delta\chi_{-}(-E)\right]\,,$ (78)
where
$\delta\chi_{\pm}(E)\equiv$ $(E_{\nu}^{\mp})^{2}f_{e}(-E)\,{\rm sgn}(E^{\mp}_{\nu})$
×$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,,$ (79)
@@
where $\Theta(x)$ is the Heaviside step function, and $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ represent the spectral distortions of electron neutrinos and antineutrinos, respectively. For a given process, the various sgn and $\Theta$ functions together characterize the nature and state of the neutrino particle, indicating whether it is a neutrino or an antineutrino and whether it is in the initial or final state. Additionally, we note that $T_{\rm CM}$ and $T_{\nu}$ exhibit slight differences after neutrino decoupling. If we ignore this difference and set $\xi_{\nu}=0$, resulting in $\delta\varrho_{\nu}=\delta\varrho_{\bar{\nu}}$, then Eq. (4.1) coincides with Eq. (29) in Ref. Froustey:2019owm, where corrections to the weak rates from neutrino spectral distortions in the SM were considered.
@@
Finally, we provide a brief overview of other corrections to the weak interaction rates considered in this work. First, we address the zero-temperature radiative corrections, which include pure radiative corrections corresponding to the emission and absorption of a virtual photon during the interaction, as well as bremsstrahlung corrections arising from the emission of a real photon. Second, we consider finite-temperature radiative corrections related to the finite temperature effects on proton-electron interactions and electron energy. Third and fourth, we include corrections due to the finite nucleon mass and weak magnetism. The total weak reaction rates, combining all these corrections, are expressed as
@@
where the superscripts ’RC0’, ’SD’, ’T’, ’BS’, and ’FM’ denote the zero-temperature radiative correction, spectral distortion, finite-temperature radiative correction, bremsstrahlung, and finite nucleon mass correction, respectively. The form for the $p\to n$ rates is analogous.
It is important to note that the effects of different corrections may be coupled. For instance, the impact of zero-temperature radiative corrections on those induced by spectral distortions is also addressed in this work. Specifically, the Born rates with zero-temperature radiative correction are given by
${\Gamma}^{{\rm RC}0}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\mathcal{F}_{+}(E)R(E,|\Delta-E|)\chi_{+}(E)+\mathcal{F}_{+}(-E)R(E,\Delta+E)\chi_{+}(-E)\right]\,,$ (81)
${\Gamma}^{{\rm RC}0}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p\left[\mathcal{F}_{-}(E)R(E,\Delta+E)\chi_{-}(E)+\mathcal{F}_{-}(-E)R(E,|E-\Delta|)\chi_{-}(-E)\right]\,.$ (82)
The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in, e.g., Ref. Pitrou:2018cgg, and are not repeated here for brevity. It is evident that the zero-temperature radiative correction modifies only the coefficients of the $\chi_{\pm}$ functions. The terms $\delta\Gamma^{\rm RC0,SD}_{n\to p}$ and $\delta\Gamma^{\rm RC0,SD}_{p\to n}$ can be derived simply by replacing $\chi_{\pm}$ with $\delta\chi_{\pm}$. In principle, other corrections are also influenced by the spectral distortions of neutrinos and antineutrinos. However, these effects are expected to be negligible, as the corrections themselves are already sub-leading. Thus, we disregard these interferences in this work.